OpenStudy (anonymous):

MEDAL! Find the center, vertices, and foci of the ellipse with equation x^2/400+y^2/625=1

OpenStudy (imstuck):

Ok, your equation looks like this:\[\frac{ x ^{2} }{ 400 }+\frac{ y ^{2} }{ 625 }=1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

center i know is (0,0)

OpenStudy (imstuck):

That is in the form \[\frac{ (x-h)^{2} }{ b ^{2} }+\frac{ (y-k)^{2} }{ a ^{2} }=1\]

OpenStudy (imstuck):

You're right when you say the center is at the origin. Good. That's the first thing out of the way.

OpenStudy (imstuck):

The second thing to remember about an ellipse is that a is ALWAYS bigger than b. So if the a (or a^2) is under the x^2, the x axis has the major vertices, and the y axis has the minor vertices.

OpenStudy (imstuck):

Our a^2 lies under our y^2 so the y axis is the major vertex and the ellipse in general will look like this:

OpenStudy (imstuck):



OpenStudy (imstuck):

See that? If the a^2 lie under the x^2, then the x axis is the major vertex and it would look like this:

OpenStudy (anonymous):

yes

OpenStudy (imstuck):



OpenStudy (anonymous):

(0, -25), (0, 25) would be the vertices right?

OpenStudy (imstuck):

So you know the major vertices lie on the y axis, but where is what we need to find.

OpenStudy (imstuck):

Wow! Yes, you'r right about the major vertices! Good job!

OpenStudy (anonymous):

yay! thanks!

OpenStudy (anonymous):

so do the foci land on the x axis?

OpenStudy (imstuck):

The formula for the foci is this:\[a ^{2}-b ^{2}=c ^{2}\]And because this is a y axis ellipse, the foci lie on the y axis as well. So filling in our formula to find the foci we have this:

OpenStudy (anonymous):

(0, -15), (0, 15) right?

OpenStudy (imstuck):

\[625-400=c ^{2}\]

OpenStudy (imstuck):

Yep, you're correct! Ellipses are pretty easy; they're a lot like circles. Good Job!

OpenStudy (anonymous):

thanks!

OpenStudy (imstuck):

The minor vertiices will lie on the x axis at (20,0) and (-20,0).

OpenStudy (imstuck):

YW! TY for the medal!

OpenStudy (anonymous):

just to be sure , this is it right?Center: (0, 0); Vertices: (0, -25), (0, 25); Foci: (0, -15), (0, 15)

OpenStudy (anonymous):

just want to make sure i did not mess up the minor and major @IMStuck